what is the slope intercept form of y+6=2(x+2)

Answer by jim_thompson5910(35256)

y+6=2(x+2)

y+6=2(x)+2(2)

y+6=2x+4

y=2x+4-6

y=2x-2


So the answer is y=2x-2